OpenStudy (chunkymonkay):

Which is an equation in point-slope form for the given point and slope? Point (2,-6) ; slope -3/4

OpenStudy (misty1212):

the point slope form is \[\huge y-\color{red}{y_1}=\color{blue}m(x-\color{green}{x_1})\]

OpenStudy (misty1212):

replace \(\color{red}{y_1}\) by \(\color{red}{6}\) and replace \(\color{\blue}m\) by \(\color{blue}{-\frac{3}{4}}\) and \(\color\green{x_1}\) by \(\color{green}2\)

OpenStudy (chunkymonkay):

so it would be y-6=3/4(x-2)

OpenStudy (chunkymonkay):

right?

OpenStudy (misty1212):

not quite

OpenStudy (misty1212):

\[\huge y-\color{red}{y_1}=\color{blue}m(x-\color{green}{x_1})\] your answer is missing the minus sign for the slope also \(y-(-6)=y+6\)

OpenStudy (misty1212):

i think i meant "you're " missing the minus sign

OpenStudy (chunkymonkay):

so then y+6=-3/4(x-2)

OpenStudy (misty1212):

yup

OpenStudy (chunkymonkay):

yayy thank you soo much

OpenStudy (misty1212):

you are quite welcome dear, happy to help